OpenStudy (anonymous):

Two systems of equations are given below. For each system, choose the best description of its solution. If applicable, give the solution. x-y=-2 -x+2y=2

OpenStudy (anonymous):

x=y-2 2y=2+x Are there descriptions to choose from?

OpenStudy (anonymous):

1.The system has no solution 2The system has a unique solution: (x,y) 3.The system has infinitely many solutions. they must satisfy the following equation y=

OpenStudy (anonymous):

@sarahusher

OpenStudy (anonymous):

Both systems have infinitely many solutions, as no matter what x or y you sub in, you can get an appropriate y or x value http://www.wolframalpha.com/input/?i=plot+x%3Dy-2%2C+2y%3D2%2Bx